OpenStudy (anonymous):

we have a continues function at [0,1] , f(0) = 0 and f(1)=1 , how do i prove that there is a number c from ]0,1[ achieve f(c) = (1-c)/(1+c) ?

OpenStudy (anonymous):

What do you think @Zarkon i know that we should use the intermediate values theorem but how ?

OpenStudy (zarkon):

look at \[g(x)=f(x)-\frac{1-x}{1+x}\]

OpenStudy (anonymous):

Great then ?

OpenStudy (zarkon):

what is g(0) and g(1)?

OpenStudy (anonymous):

g(0) = f(0) - 1 ?

OpenStudy (zarkon):

yes =-1

OpenStudy (anonymous):

g(1) = f(1)

OpenStudy (zarkon):

=1

OpenStudy (anonymous):

exactly

OpenStudy (zarkon):

you have it from here?

OpenStudy (anonymous):

then ?

OpenStudy (zarkon):

use the IVT

OpenStudy (anonymous):

for g or f ?

OpenStudy (zarkon):

g

OpenStudy (zarkon):

g(0)<0 and g(1)>0 so...

OpenStudy (anonymous):

g(0)*g(1) < 0

OpenStudy (anonymous):

so there is a c for g , what about f ?

OpenStudy (zarkon):

got there is a c such that g(c)=0 so \[f(c)-\frac{1-c}{1+c}=0\]

OpenStudy (anonymous):

mmmm cool

OpenStudy (anonymous):

@Zarkon how did you think about the g(x) function , is there a princible ?